b) Length of a rectangle is 3 more than its width. The perimeter of the rectangle is 30. Find the length and width.

Found 2 solutions by algebrahouse.com, mananth:
Answer by algebrahouse.com(1659)

"b) Length of a rectangle is 3 more than its width. The perimeter of the rectangle is 30. Find the length and width."

x = width
x + 3 = length {length is 3 more than width}

Perimeter of a rectangle = 2(width) + 2(length}

2x + 2(x + 3) = 30 {perimeter is 2(width) + 2(length)}
2x + 2x + 6 = 30 {used distributive property}
4x + 6 = 30 {combined like terms}
4x = 24 {subtracted 6 from both sides}
x = 6 {divided both sides by 4}
x + 3 = 9 {substituted 6, in for x, into x + 3}

width = 6 and length = 9
www.algebrahouse.com

Answer by mananth(16946)

Perimeter = 30
Length =x+3
Width =xcm

Perimeter =2((x +3)+x)
Perimeter =2(2x+ 3)
Perimeter =4x +6
30=4x+6
30-6=4x
24=4x
x=6 Width
Length =9